# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v4

@@ 5. Numerical Results
For more details about TFI system, we refer the reader to Appendix C. For the TFI system, the cp-AFQMC can be made free of sign problem, as analyzed in detail in Appendix C. Without a sign problem, the constrained-path approximation will not incur any systematic error, and the trial wavefunction $\Psi_{\mathrm{tr}}$ purely introduces importance sampling. Depending on the quality of $\Psi_{\mathrm{tr}}$, the efficiency as well as other controllable biases in the computation are affected, including those from fixed time step $\Delta\tau$ and finite population size $N$, which are discussed further in Appendix D. These provide an excellent test ground to illustrate and study our algorithm. As we show below, the algorithm leads to significant improvement across the board by allowing better importance sampling. Furthermore, by combining the features of AFQMC and DMRG, the algorithm shows advantages and extends the reach beyond each of its components separately.
@@
where $\Psi_{0}(g)$ represents the ground-state wavefunction of the system with the transverse-field strength being $g$. Numerically, this derivative is approximated by finite difference:
@@
Since the quantum phase transition at the critical point $g=1$ makes it the most difficult case to deal with, we also compare cp-AFQMC-re-anchoring and the vanilla cp-AFQMC on larger systems, including 1D system with $d=32,64$ and $96$ spins, at $g=1$. The ground-state energy of 1D periodic system can be calculated analytically [50], and the ground-state wavefunction can be well approximated by density matrix renormalization group (DMRG) [51, 52].
@@
Figure 5. The overlap between the ground-state wavefunction and the sketched wavefunction from cp-AFQMC-re-anchoring, in the form of a rank-$4$ TT. From (a) to (d): $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
@@
Table 1. Relative error of the energy. The new algorithm leads to a reduction of statistical error by more than two orders of magnitude, or $\mathcal{O}(10,000)$ reduction in computer time. Systematic errors are also reduced. The systematic errors still present in cp-AFQMC comes from the fixed imaginary-time step and the finite population size, which can be eliminated by extrapolating. See Appendix D for detailed discussion.
As an example of a 2D system, we test our algorithm on the $11\times 1$ 11 lattice with open boundary at the quantum critical point $g\approx 3.044$ [53]. In our experiments, we take $g=3$. The results are shown in Figure 6. For such a 2D system, there is no analytic solution for the ground-state energy anymore. A reference for the ground-state energy given by DMRG is $-3.17210\pm 1\times 10^{-5}$ [54], while the energy estimated by our method is $-3.17212\pm 4\times 10^{-5}$.
@@
Figure 6. Energy convergence of (a) vanilla cp-AFQMC and (b) cp-AFQMC-re-anchoring for $11\times 1$ 11 system with open boundary at $g=3$. The reference ground-state energy, shown as the black dashed line, is obtained from DMRG [54].
@@ 5.2. Comparison with DMRG
Figure 7. Comparison between cp-AFQMC-re-anchoring and DMRG for the $4\times 4$ system with periodic boundary with different transverse-field strengths. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by DMRG and TT-sketching; (b): comparison of the relative error in ground-state energy estimation by cp-AFQMC-DMRG and cp-AFQMC-re-anchoring.
Consistent with the observation above, the computed ground-state energies also show an improvement with the algorithm, as illustrated in Figure 7(b). We take a closer look at $g=2.0$, where the two algorithms exhibit a significant difference. The overlap between the ground-state and the sketched wavefunction during the cp-AFQMC-re-anchoring procedure is shown in Figure 8(a). In terms of the ground-state energy, Figure 8(b) clearly shows that cp-AFQMC-re-anchoring provides a more accurate estimation compared to cp-AFQMC-DMRG. Specifically, the relative energy error for cp-AFQMC-re-anchoring is $(0.59\pm 1.57)\times 10^{-5}$, while the error for cp-AFQMC-DMRG is $(+1.87\pm 0.08)\times 10^{-4}$, which exhibits a substantial bias.
@@
Figure 8. Comparison between cp-AFQMC-re-anchoring and cp-AFQMC-DMRG for the $4\times 4$ system with periodic boundary at $g=2.0$. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by both methods, in the form of a rank-$4$ TT; (b): comparison of the energy convergence.
@@
where $\braket{O}_{\mathrm{mix}}$ denotes the mixed estimator, analogous to (13), and
@@
(Because the quality of our trial wavefunction is very high, the extrapolated estimator is sufficient to remove any residual systematic error.) As an example, we study the average spin in $z$-direction:
@@
Due to the symmetry of the Hamiltonian (27) under spin-flip transformations in $\sigma_{z}$, specifically the invariance of the Hamiltonian when all $\sigma_{i}^{z}$ are flipped to $-\sigma_{i}^{z}$, the expectation value of $\mathcal{S}_{z}$ in the ground state is $\braket{\mathcal{S}_{z}}_{\mathrm{gs}}=0$. Numerical results in Figure 9 demonstrate that cp-AFQMC-re-anchoring preserves the $\sigma_{z}$ symmetry, which is broken in cp-AFQMC-DMRG.
@@
Figure 10. Spin-correlation (defined in (30)) of the $4\times 4$ system with periodic boundary at $g=2.0$. (a): labeling of the spins; (b): correlations with the reference spin $1$.
@@ Acknowledgments
YK is partially funded by DMS-2339439, DE-SC0022232, and Sloan Research Fellowship.
@@ Appendix A Time Complexity of TT Calculation
Figure 13. Computational cost of TT-sketching. (a): different number of spins with $2000$ walkers; (b): different number of walkers for $16$ spins.
@@ Appendix B Details on Convergence Analysis of cp-AFQMC
(37) $\frac{\braket{\Psi_{\mathrm{tr}},H(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(38) $=$ $\frac{\braket{H\Psi_{\mathrm{tr}},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(39) $=$ $\frac{\braket{c_{0}E_{0}\Psi_{0}+c_{\perp}H\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(40) $=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(41) $=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
(42) $\leq$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
(43) $=$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|c_{0}\braket{\Psi_{0},e^{-\Delta\tau H}\Phi_{k}^{(n)}}+c_{\perp}\braket{\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
(44) $\leq$ $\|H-E_{0}I\|_{2}\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\left|\frac{c_{\perp}}{c_{0}}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{\braket{\Psi_{0},\Phi_{k}^{(n)}}}\right|\bigg{/}\left(1+a_{0}e^{-\Delta(E_{1}-E_{0})}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$
(45) $=$ $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\frac{a_{0}\tan\theta_{k}^{(n)}\|H-E_{0}I\|_{2}}{1+a_{0}e^{-\Delta(E_{1}-E_{0})}\tan\theta_{k}^{(n)}},$
where second equality follows from (22) and Cauchy–Schwarz inequality, and we use $\mathcal{P}_{\perp}$ to denote the orthogonal projector that projects to the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. Plugging this and the bound in Lemma 4 into (36) then gives (26). ∎
@@
$\begin{split}&\quad\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=\braket{c_{0}\Psi_{0}+c_{\perp}\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=c_{0}\braket{e^{-\Delta\tau E_{0}}\Psi_{0},\Phi_{k}^{(n)}}+c_{\perp}\braket{\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}},\\ \end{split}$
@@
$\begin{split}\braket{\Psi_{0},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=e^{-\Delta\tau E_{0}}\braket{\Psi_{0},\Phi_{k}^{(n)}}-\frac{c_{\perp}}{c_{0}}\braket{\Psi_{\perp},E^{(n)}_{k}}.\end{split}$
@@
$\begin{split}\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}\Phi_{k}^{(n)}\|&\leq\|\mathcal{P}_{\perp}e^{-\Delta\tau H}\Phi_{k}^{(n)}\|+\|\mathcal{P}_{\perp}\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|\leq e^{-\Delta\tau E_{1}}\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|+\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|,\\ \end{split}$
@@
Appendix C cp-AFQMC on Transverse-Field Ising Model
In this section, we follow the procedure of the CPMC algorithm on Hubbard model [57] to solve for TFI model. For the Hamiltonian (27), we use the Suzuki-Trotter approximation [58, 59] to approximate $e^{-\Delta\tau H}$:

### v5

@@ 5. Numerical Results
For more details about TFI system, we refer the reader to the appendix. For the TFI system, the cp-AFQMC can be made free of sign problem; a detailed analysis is provided in the supplementary appendix. Without a sign problem, the constrained-path approximation will not incur any systematic error, and the trial wavefunction $\Psi_{\mathrm{tr}}$ purely introduces importance sampling. Depending on the quality of $\Psi_{\mathrm{tr}}$, the efficiency as well as other controllable biases in the computation are affected, including those from fixed time step $\Delta\tau$ and finite population size $N$. See further discussions in the appendix. These provide an excellent test ground to illustrate and study our algorithm. As we show below, the algorithm leads to significant improvement across the board by allowing better importance sampling. Furthermore, by combining the features of AFQMC and the density matrix renormalization group (DMRG) [52, 15], the algorithm shows advantages and extends the reach beyond each of its components separately.
@@
where $\Psi_{0}(g)$ represents the ground-state wavefunction of the TFI system with the transverse-field strength being $g$. Numerically, this derivative is approximated by finite difference:
@@
Since the quantum phase transition at the critical point $g=1$ makes it the most difficult case to deal with, we also compare cp-AFQMC-re-anchoring and the vanilla cp-AFQMC on larger systems, including 1D system with $d=32,64$ and $96$ spins, at $g=1$. The ground-state energy of 1D periodic system can be calculated analytically [29], and the ground-state wavefunction can be well approximated by DMRG.
@@
Figure 6. The overlap between the ground-state wavefunction and the sketched wavefunction from cp-AFQMC-re-anchoring, in the form of a rank-$4$ $T$. From (a) to (d): $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
@@
Table 1. Relative error of the energy. The new algorithm leads to a reduction of statistical error by more than two orders of magnitude, or $\mathcal{O}(10,000)$ reduction in computer time. Systematic errors are also reduced. The systematic errors still present in cp-AFQMC comes from the fixed imaginary-time step and the finite population size, which can be eliminated by extrapolating. A detailed discussion is provided in the appendix.
As an example of a 2D system, we test our algorithm on the $11\times 11$ lattice with open boundary at the quantum critical point $g\approx 3.044$ [6]. In our experiments, we take $g=3$. The results are shown in Figure 7. For such a 2D system, there is no analytic solution for the ground-state energy anymore. A reference for the ground-state energy given by DMRG is $3.17210\pm 1\times 10^{-5}$ [23], while the energy estimated by our method is $3.17212\pm 4\times 10^{-5}$.
@@
Figure 7. Energy convergence of (a) vanilla cp-AFQMC and (b) cp-AFQMC-re-anchoring for $11\times 11$ system with open boundary at $g=3$. The reference ground-state energy, shown as the black dashed line, is obtained from DMRG [23].
@@ 5.2. Comparison with DMRG
Figure 8. Comparison between cp-AFQMC-re-anchoring and DMRG for the $4\times 4$ system with periodic boundary with different transverse-field strengths. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by DMRG and $T$-sketching; (b): comparison of the relative error in ground-state energy estimation by cp-AFQMC-DMRG and cp-AFQMC-re-anchoring.
In addition to the fact that our sketched wavefunction generally provides a better approximation of the ground state, cp-AFQMC-re-anchoring can also yield a more accurate estimation of the ground-state energy than cp-AFQMC with DMRG wavefunction as the trial function (cp-AFQMC-DMRG). We take a closer look at $g=2.0$, where the two algorithms exhibit a significant difference. The overlap between the ground-state and the sketched wavefunction during the cp-AFQMC-re-anchoring procedure is shown in Figure 9(a). In terms of the ground-state energy, Figure 9(b) clearly shows that cp-AFQMC-re-anchoring provides a more accurate estimation compared to cp-AFQMC-DMRG. Specifically, the relative energy error for cp-AFQMC-re-anchoring is $(0.59\pm 1.57)\times 10^{-5}$, while the error for cp-AFQMC-DMRG is $(+1.87\pm 0.08)\times 10^{-4}$, which exhibits a substantial bias.
@@
Figure 9. Comparison between cp-AFQMC-re-anchoring and cp-AFQMC-DMRG for the $4\times 4$ system with periodic boundary at $g=2.0$. (a): comparison of the overlap between the ground-state wavefunction and the wavefunctions obtained by both methods, in the form of a rank-$4$ $T$; (b): comparison of the energy convergence.
@@
where $\braket{O}_{\mathrm{mix}}$ denotes the mixed estimator
$\braket{O}_{\mathrm{mix}}=\frac{\braket{\Psi_{\mathrm{tr}},O\widehat{\Psi}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widehat{\Psi}^{(n)}}},$
which is estimated from the cp-AFMQC procedure, and
@@
Because the quality of our trial wavefunction is very high, the extrapolated estimator is sufficient to remove any residual systematic error. As an example, we study the average spin in $z$-direction:
@@
Due to the symmetry of the Hamiltonian (20) under spin-flip transformations in $\sigma_{z}$, specifically the invariance of the Hamiltonian when all $\sigma_{i}^{z}$ are flipped to $\sigma_{i}^{z}$, the expectation value of $\mathcal{S}_{z}$ in the ground state is $\braket{\mathcal{S}_{z}}_{\mathrm{gs}}=0$. Numerical results in Figure 10 demonstrate that cp-AFQMC-re-anchoring preserves the $\sigma_{z}$ symmetry, which is broken in cp-AFQMC-DMRG.
@@
Figure 11. Spin-correlation (defined in (23)) of the $4\times 4$ system with periodic boundary at $g=2.0$. (a): labeling of the spins; (b): correlations with the reference spin 11.
@@ Acknowledgments
YK is partially funded by DMS-2339439, DE-SC0022232, and Sloan Research Fellowship. The Flatiron Institute is a division of the Simons Foundation.
@@ Appendix A Time Complexity of TT Calculation
Figure 14. Computational cost of $T$-sketching. (a): different number of spins with $2000$ walkers; (b): different number of walkers for $16$ spins.
@@ Appendix B Details on Convergence Analysis of cp-AFQMC
$\frac{\braket{\Psi_{\mathrm{tr}},H(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
$=$ $\frac{\braket{H\Psi_{\mathrm{tr}},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
$=$ $\frac{\braket{c_{0}E_{0}\Psi_{0}+c_{\perp}H\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
$=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
$=$ $\frac{\braket{c_{\perp}(H-E_{0}I)\Psi_{\perp},(\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H})\Phi_{k}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}}$
$\leq$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
$=$ $c_{\perp}\|H-E_{0}I\|_{2}\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|c_{0}\braket{\Psi_{0},e^{-\Delta\tau H}\Phi_{k}^{(n)}}+c_{\perp}\braket{\Psi_{\perp},e^{-\Delta\tau H}\Phi_{k}^{(n)}}|}$
$\leq$ $\|H-E_{0}I\|_{2}\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\left|\frac{c_{\perp}}{c_{0}}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{\braket{\Psi_{0},\Phi_{k}^{(n)}}}\right|\bigg/\left(1+a_{0}e^{-\Delta(E_{1}-E_{0})}\frac{\|\mathcal{P}_{\perp}\Phi_{k}^{(n)}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$
$=$ $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|_{2}}{e^{-\Delta\tau E_{0}}}\frac{a_{0}\tan\theta_{k}^{(n)}\|H-E_{0}I\|_{2}}{1+a_{0}e^{-\Delta(E_{1}-E_{0})}\tan\theta_{k}^{(n)}},$
where the second equality follows from (15) and Cauchy–Schwarz inequality, and we use $\mathcal{P}_{\perp}$ to denote the orthogonal projector onto the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. Plugging this and the bound in Lemma 4 into (29) then gives (19). ∎
@@
[…]
@@
[…]
@@
[…]
@@
Appendix C Implementation of cp-AFQMC on Transverse-Field Ising Model
In this section, we introduce the detailed implementation of Algorithm 1, and how we apply it on the transverse-field Ising model.
C.1. Implementation details of cp-AFQMC
To implement Algorithm 1, we rewrite the wavefunction representation (3) as
(31) $\widehat{\Psi}^{(n)}=\sum_{k}w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}},$
where $w_{k}^{(n)}$ denotes the weight of the $k$-th walker. In actual implementations, a walker is fully specified by $\phi_{k}^{(n)}$, its associated weight $w_{k}^{(n)}$, and the overlap $\langle\Psi_{\mathrm{tr}},\phi_{k}^{(n)}\rangle$, which are stored and updated during the simulation. In particular, whenever $\phi_{k}^{(n)}$ is updated, the overlap $\langle\Psi_{\mathrm{tr}},\phi_{k}^{(n)}\rangle$ is updated accordingly. Since the random walkers are guaranteed to always have a positive overlap with the trial wavefunction in cp-AFQMC, the weight of a valid walker is always positive $w_{k}^{(n)}>0$. The propagation of the walkers (8) can then be rewritten as
$e^{-\Delta\tau H}\widehat{\Psi}^{(n)}$ $\approx$ $\sum_{k}\sum_{\vec{x}:Q^{(n)}_{k}(\vec{x})>0}Q^{(n)}_{k}(\vec{x})w^{(n)}_{k}\frac{P(\vec{x})}{Q^{(n)}_{k}(\vec{x})}\frac{B(\vec{x}){\phi}_{k}^{(n)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}}$
$\approx$ $\sum_{k}w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{Q^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{\phi_{k}^{(n+1)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}}$
$=$ $\sum_{k}\left(w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{Q^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n+1)}}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}}\right)\frac{\phi_{k}^{(n+1)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n+1)}}}$
$=$ $\sum_{k}w_{k}^{(n+1)}\frac{\phi_{k}^{(n+1)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n+1)}}},$
where
$\phi_{k}^{(n+1)}=B(\vec{x}^{(n)}_{k}){\phi}_{k}^{(n)},\vec{x}^{(n)}_{k}\sim Q^{(n)}_{k}$,
and
(32) $w^{(n+1)}_{k}=w^{(n)}_{k}\mathcal{N}_{k}^{(n)}.$
With the wave function representation (31), we can also estimate the ground-state energy in a straightforward way. Specifically, we use the mixed estimator
where $E_{\text{local}}$ represents the local energy of a walker:
As the random walk goes on, the weights of the walkers can become highly uneven: some grow very large while others shrink to negligible values. Consequently, only a small fraction of walkers effectively contribute to the energy estimation, leading to increased variance and reduced sampling efficiency. To solve this issue, we periodically apply a population control procedure, in which walkers with very small weights are removed, while those with large weights are branched into multiple copies. The detailed implementation is summarized in Algorithm 3.
Algorithm 3 Population control
walkers $\phi_{k}$, weights $w_{k}$, and corresponding $\braket{\Psi_{\mathrm{tr}},\phi_{k}}$.
adjusted walkers $\phi_{k}$, weights $w_{k}$, and corresponding $\braket{\Psi_{\mathrm{tr}},\phi_{k}}$.
normalize the weights so that the average value of the weights is $1$.
$s\leftarrow\text{a uniformly random number in }(0,1)$,
$n_{1}\leftarrow 1$,
for $k=1,2,\cdots,N$ do
  $s\leftarrow s+w_{k}$;
[…]
  set the walkers $\phi_{n_{1}},\cdots,\phi_{n_{2}-1}$ to be $\phi_{k}$, and corresponding overlaps to be $\braket{\Psi_{\mathrm{tr}},\phi_{k}}$;
  $n_{1}\leftarrow n_{2}$
end for
reset all the weights to be $w_{k}=1$.
C.2. cp-AFQMC on TFI model
To apply the cp-AFQMC procedure on TFI model, we follow the procedure of the CPMC algorithm on Hubbard model [26]. For the Hamiltonian (20), we use the Suzuki-Trotter approximation [44, 45] to approximate $e^{-\Delta\tau H}$: